What's the rule of (0,-1),(1,0),(2,7),(3,26),(4,63)

Answer by Theo(13342)

i believe it's y = x^3 - 1

0^3 - 1 = -1
1^3 - 1 = 0
2^3 - 1 = 7
3^3 - 1 = 26
4^3 - 1 = 63